### v1

# GIM: Improved Interpretability for Large Language Models

Abstract
Ensuring faithful interpretability in large language models is imperative for trustworthy and reliable AI. A key obstacle is self-repair, a phenomenon where networks compensate for reduced signal in one component by amplifying others, masking the true importance of the ablated component. While prior work attributes self-repair to layer normalization and back-up components that compensate for ablated components, we identify a novel form occurring within the attention mechanism, where softmax redistribution conceals the influence of important attention scores. This leads traditional ablation and gradient-based methods to underestimate the significance of all components contributing to these attention scores. We introduce Gradient Interaction Modifications (GIM), a technique that accounts for self-repair during backpropagation. Extensive experiments across multiple large language models (Gemma 2B/9B, LLAMA 1B/3B/8B, Qwen 1.5B/3B) and diverse tasks demonstrate that GIM significantly improves faithfulness over existing circuit identification and feature attribution methods. Our work is a significant step toward better understanding the inner mechanisms of LLMs, which is crucial for improving them and ensuring their safety. Our code is available at https://github.com/JoakimEdin/gim.
Figure 1: Attention self-repair. When multiple values associated with large attention weights contain similar information, ablating one attention score has little effect on the output because the softmax activation function compensates by increasing the weight of other positions. This results in perturbation-based and gradient-based explanation methods underestimating the importance of components that contribute to the attention scores through the keys and queries.
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, from text summarization to code synthesis [1]. Despite their impressive performance, our understanding of their inner mechanisms remains limited. This opacity hinders our ability to explain model outputs, catch hallucinations, systematically improve performance, and ensure reliable behavior in novel situations. To address this, most of the previous explainability methods aim to quantify how internal components (e.g., attention heads, neurons, token embeddings) affect model output [2]. Two of the most widely adopted approaches, perturbation-based and gradient-based methods, both rest on the assumption that modifying a truly salient component will lead to a commensurate change in the model’s output. For instance, perturbation-based methods (e.g., SHAP, LIME, activation patching) investigate the effect of large perturbations [3, 4], while gradient-based methods analyze infinitesimally small perturbations [5]. However, this core assumption that perturbing important components will perturb the model output often fails due to what is known as the self-repair effect [6].
Self-repair is a phenomenon in language models where downstream model components compensate for perturbed components, resulting in deceptively small output changes [7, 8, 6]. As a consequence, explanation methods may underestimate the importance of certain model components [6].
In this paper, we identify a new form of self-repair within the attention mechanism of LLMs. When multiple value vectors associated with large attention weights contain similar information, perturbing any individual attention score has little effect on the model output. If one position’s attention score is reduced, the softmax activation function increases the attention weight of the other positions, maintaining similar output signals (see Figure 1). This attention self-repair systematically causes both perturbation-based methods and gradient methods to underestimate the importance of all components contributing to these attention scores.
To address these challenges, we propose gradient interaction modifications (GIM), a novel gradient-based attribution method specifically designed to account for self-repair effects in language models. GIM introduces targeted modifications to backpropagation to better handle interactions that cause self-repair. First, we develop temperature-adjusted softmax gradients to address the attention self-repair problem. Second, we implement layernorm freeze to deal with the self-repair in layer normalization [9, 6]. Finally, we incorporate gradient normalization as proposed by Achtibat et al. [10], which we found essential for making our backpropagation modifications effective in practice.
Our key contributions are:
  1. 1.
    We identify and formalize the attention self-repair problem in softmax operations, showing how it fundamentally undermines the faithfulness metrics of traditional interpretability methods.
  2. 2.
    We introduce GIM, a novel gradient-based attribution method that combines three complementary techniques to address feature interactions in transformer components. We demonstrate that GIM is more faithful than other gradient-based circuit identification and feature attribution methods across multiple large language models (Gemma 2B/9B, LLAMA 1B/3B/8B, and Qwen 1.5B/3B) and six datasets spanning four tasks (question-answering, fact-verification, sentiment classification, and hate speech detection).
  3. 3.
    We empirically isolate and demonstrate the importance of each of our three modifications, showing their significant positive effect on explanation faithfulness across multiple large language models and tasks.
2 Related Work
2.1 Self-repair in Neural Networks
Self-repair occurs when ablating components have unexpectedly small effects on model outputs because other components compensate for the ablation. Several studies have identified attention heads that suppress the signal from earlier layers, and the suppression is reduced when those early layers are ablated  [8, 7, 11]. McGrath et al. [8] identified what they called MLP erasure neurons that have a similar suppressing effect on the signal from previous layers. Rushing and Nanda [6] investigated the role of layer normalization in self-repair. They demonstrated that normalization can compensate for ablated layers by rescaling the remaining signals. Our work identifies a new form of self-repair that occurs within attention mechanisms, where softmax redistribution conceals the influence of important attention scores. We also show that this self-repair mechanism can impact the gradient.
2.2 Gradient-based Attribution Methods
Attribution methods quantify model components’ contribution to specific model behaviors. Feature attribution methods quantify the importance of input features (e.g., tokens), while circuit identification methods quantify the importance of internal model components (e.g., attention heads). Attribution methods are usually perturbation-based or gradient-based [12]. Perturbation-based approaches (e.g., SHAP [3] and activation patching [13]) attribute importance by measuring output changes when perturbing components. However, they require numerous forward passes, making them prohibitively slow for large language models. Gradient-based methods [5, 14] offer a faster alternative by leveraging backpropagation. However, they often produce unfaithful explanations due to neural network non-linearities, where first-order approximations fail to capture true component importance [10, 15]. Methods such as Layer-wise Relevance Propagation (LRP) [16], Integrated Gradients [5], and DeepLIFT [17] attempt to address these limitations by modifying gradient flows, but empirically struggle with language models [18, 19, 10].
Several studies propose specialized backpropagation rules for the transformer architecture. TransformerLRP [9] treats the layer normalization scaling variable and the attention matrix as constants, essentially not backpropagating through them. ATP* [20] propose a backpropagation rule for the softmax operation that better estimates the impact of large attention score perturbations. However, ATP* ignores the normalization contribution in the softmax operation. Achtibat et al. [10] propose AttnLRP, which is a collection of backpropagation rules. Grad norm, their most important contribution to our work, divides the gradient by $N$ when backpropagating through operations where $N$ variables are multiplied together. [^1] Achtibat et al. refer to grad norm as the uniform rule Using Taylor Decomposition and Shapley, they provide theoretical justification for this rule. While TransformerLRP and AttnLRP show that the combination of their backpropagation rules improves the attribution faithfulness, they do not isolate each rule’s contribution to the improvements.
Unlike previous work, our approach specifically targets self-repair in attention mechanisms and systematically evaluates both existing and novel backpropagation modifications, measuring their individual and combined impacts on attribution faithfulness.
3 Attention Self-Repair
Self-repair in the attention layer is when an attention score strongly influences the output, yet perturbing it has minimal effect. This effect deems both perturbation- and gradient-based attribution methods inaccurate. In this section, we show when and why attention self-repair occurs and its consequences on gradient-based explanation methods.
In transformer models, attention mechanisms control information flow between positions. Position $i$ gathers information from all positions using the following computations:
$\begin{split}&\bm{a}=\bm{Q}_{i}\cdot\bm{K}^{T}\quad\text{(Attention scores for position $i$)}\\ &\bm{s}=\text{Softmax}(\bm{a})\quad\text{(Attention weights for position $i$)}\\ &\bm{o}=\bm{s}\cdot\bm{V}\quad\text{(Attention output for position $i$)}\end{split}$ (1)
where $\bm{Q}_{i}$ is the query for position $i$, $\bm{K}$ are the keys, and $\bm{V}$ are the values. The softmax function normalizes attention scores into a probability distribution:
$\text{Softmax}(\bm{a})=\frac{e^{\bm{a}/\tau}}{\sum_{k}e^{\bm{a}_{k}/\tau}}$ (2)
where $\tau$ is the temperature parameter. For clarity, we distinguish between softmax input “attention scores” $\bm{a}$ and softmax output “attention weights” $\bm{s}$.
Attention mechanisms can be viewed as information routing systems, where each value vector $\bm{V}_{j}$ contain information from position $j$, and the attention weights $\bm{s}_{j}$ determine how much information to copy from that position.
Self-repair occurs when multiple values with large attention weights contain similar information. When an attention score $\bm{a}_{j}$ is ablated, the softmax redistributes weights primarily to positions with the highest remaining scores. If these positions contain values $\bm{V}_{j}$ that contain similar information, the output remains virtually unchanged.
Figure 1 illustrates this with attention weights $[0.5,0.5,0]$ and value vectors $[1,1,0]$. Note that the values are identical at the positions with non-zero attention weights. Individually ablating either attention score (i.e., softmax input) leaves the output unchanged, as the softmax shifts weight to the other position with the same value. Only ablating both scores simultaneously affects the output, as the softmax shifts the attention weights to the position with a different value.
3.1 Attention self-repair results in zero gradient
Here we will prove that when the conditions for attention self-repair are met, the gradient of the model’s final output logit $z$ with respect to the attention score $\bm{a}_{j}$ will be near-zero. We can define the gradient as:
$\frac{\partial z}{\partial\bm{a}_{j}}=\bm{s}_{j}\left(\frac{\partial z}{\partial\bm{s}_{j}}(1-\bm{s}_{j})-\sum_{k\neq j}\frac{\partial z}{\partial\bm{s}_{k}}\bm{s}_{k}\right)$ (3)
where $\frac{\partial z}{\partial\bm{s}_{j}}$ is the gradient of the output logit with respect to the attention weight $\bm{s}_{j}$. This is computed as follows:
$\frac{\partial z}{\partial\bm{s}_{j}}=\frac{\partial z}{\partial\bm{o}}\cdot\bm{V}_{j}$ (4)
To demonstrate the effect of attention self-repair on Equation 3, we first define a set of positions with substantial attention weights $\mathcal{I}_{\epsilon}=\{k:\bm{s}_{k}>\epsilon\}$ where $\epsilon$ is close to zero. Recall that for self-repair to occur, the values $V_{j}$ at the positions with substantial attention weights must contribute similarly to the output. We can measure a value’s unweighted contribution to the output with the dot product between the value and its upstream gradient (GradientXInput). This is equivalent to $\frac{\partial z}{\partial\bm{s}_{j}}$ in Equation 4.
Therefore, the key condition for self-repair, which occurs when values at positions with substantial attention weights have similar contributions to the output, can be expressed mathematically as:
$\frac{\partial z}{\partial\bm{s}_{j}}\approx\frac{\partial z}{\partial\bm{s}_{k}}\approx c\quad\forall j,k\in\mathcal{I}_{\epsilon}$ (5)
where $c$ is some constant. Next, we use the property that softmax weights sum to 1:
$\sum_{k\in\mathcal{I}_{\epsilon}}\bm{s}_{k}\approx 1\Leftrightarrow 1-\bm{s}_{j}\approx\sum_{\begin{subarray}{c}k\in\mathcal{I}_{\epsilon}\\ k\neq j\end{subarray}}\bm{s}_{k}$ (6)
Substituting these into Equation 3, we get:
$\begin{split}\frac{\partial z}{\partial\bm{a}_{j}}&\approx\bm{s}_{j}\bigl{(}c\sum_{\begin{subarray}{c}k\in\mathcal{I}_{\epsilon}\\ k\neq j\end{subarray}}\bm{s}_{k}-\sum_{\begin{subarray}{c}k\in\mathcal{I}_{\epsilon}\\ k\neq j\end{subarray}}c\bm{s}_{k}\bigr{)}=\bm{s}_{j}c\bigl{(}\sum_{\begin{subarray}{c}k\in\mathcal{I}_{\epsilon}\\ k\neq j\end{subarray}}\bm{s}_{k}-\sum_{\begin{subarray}{c}k\in\mathcal{I}_{\epsilon}\\ k\neq j\end{subarray}}\bm{s}_{k}\bigr{)}=0\end{split}$ (7)
This result reveals that standard gradient-based explanation methods underestimate the importance of attention scores when attention self-repair conditions are met. Even when an attention score $\bm{a}_{j}$ significantly impacts the model’s prediction, its gradient can be near-zero. This mathematical insight motivates our approach to modify gradient calculations to better reflect the true importance of components.
4 Methods
To address self-repair and other gradient distortion issues in transformer models, we introduce gradient interaction modification (GIM), a comprehensive approach combining three complementary modifications of standard back-propagation. This section describes the three modifications: Temperature-adjusted softmax gradients (TSG), Layernorm freeze, and Grad norm.
4.1 Temperature-adjusted softmax gradients
TSG modifies backpropagation through attention to address the self-repair problem presented in Section 3.1. The goal of TSG is to preserve attribution magnitude when standard backpropagation would otherwise have underestimated the importance of attention scores because of the self-repair effect. TSG addresses the issue with a simple approach:
  1. 1.
    During the forward pass, TSG uses a standard temperature ($T=1$), preserving the model’s normal behavior and predictions
  2. 2.
    During the backward pass, TSG recomputes the softmax with a higher temperature ($T>1$) before computing gradients
Increasing the temperature results in a more uniform attention weight distribution (an increase of entropy). In other words, large attention weights decrease while small weights increase. This increases our set of positions with significant weights $\mathcal{I}_{\epsilon}=\{k:\bm{s}_{k}>\epsilon\}$. Recall that for self-repair to occur, the gradients $\frac{\partial z}{\partial\bm{s}_{j}}$ must be approximately uniform across all positions in $\mathcal{I}_{\epsilon}$. By expanding $\mathcal{I}_{\epsilon}$ to include more positions, we potentially disrupt the uniformity condition that leads to cancellation. When this uniformity condition is broken, the two terms in Equation 3 may no longer completely cancel each other out, allowing for non-zero gradients.
4.2 Layernorm freeze & gradient norm
Layernorm freeze addresses the self-repair effect in the layer normalization by treating the normalization parameters as constants during backpropagation [6, 9]. Specifically, during the forward pass, we save the normalization factor $\sigma$, and during the backward pass, we divide the upstream gradient by the normalization factor $\sigma$ instead of backpropagating through it. This naively prevents the self-repair effect by preventing layer normalization from changing due to perturbations.
Gradient norm normalizes the gradient after variable multiplications [10]. During backpropagation, it divides the gradient by the number of inputs involved in the multiplication. Achtibat et al. [10] proved the optimality of this approach using Taylor decomposition and Shapley values. Specifically, we divide the gradient by 2 for three key multiplicative interactions in transformer models: 1) attention-value, 2) query-key, and 3) MLP gate-projection.
(a) Difference between ablating the two largest attention scores separately (x-axis) and jointly (y-axis). The 95% quantile interaction is colored orange.
(b) Comparison of gradient and TSG. The attention vectors with the 95% quantile interactions are colored orange.
Figure 2: The attention self-repair effect and how temperature-adjusted softmax gradients increase the gradient when there is self-repair for LLAMA-3.2 1B on FEVER. The left plot compares ablating the two largest attention scores separately and jointly. The strongest interactions are colored orange. The right plot compares the gradient for the same attention scores with the TSG. TSG results in a stronger signal for attention scores with a strong interaction.
5 Experimental setup
Our experiments were designed around three questions: 1) Does attention self-repair occur in large language models? 2) Does GIM produce more faithful explanations than existing circuit identification and feature attribution methods? 3) How much does each of GIM’s three modifications (TSG, layernorm freeze, and grad norm) contribute to the increased faithfulness? This section describes how we ran the experiments.
5.1 Models, datasets, and evaluation metrics
For our experiments, we use seven instruction-tuned large language models spanning three model families of various sizes: Gemma-2 (2B, 9B), LLaMA-3.2 (1B, 3B), LLaMA-3.1 (8B), and Qwen-2.5 (1.5B, 3B) [21, 22, 23].
We use six datasets spanning four tasks: BoolQ (question-answering), FEVER and SciFact (fact verification), Twitter sentiment classification and Movie review (sentiment classification), and HateXplain (hate speech detection). These datasets were selected based on their diversity and prevalence in prior attribution studies [24, 25, 12]. We provide more details on the datasets in Section 9.2.
We evaluate the faithfulness of the attributions using Comprehensiveness and Sufficiency [24]. Comprehensiveness ($\uparrow$) measures the average output change after cumulatively ablating top-attributed features:
$\text{Comp}(f,\bm{x},\bm{e})=\frac{1}{Nf(\bm{x})}\sum_{i=1}^{N}f(\bm{x})-f(p(\bm{x},\text{rank}(\bm{e})_{1:i}))$ (8)
where $f$ is the model, $\bm{x}$ is the input, $\bm{e}$ contains attribution scores, $p$ is the perturbation function, and $\text{rank}(\cdot)$ orders features by attribution.
Sufficiency ($\downarrow$) measures the average output change after cumulatively ablating the lowest-attributed features:
$\text{Suff}(f,\bm{x},\bm{e})=\frac{1}{Nf(\bm{x})}\sum_{i=1}^{N}f(\bm{x})-f(p(\bm{x},\text{rank}(-\bm{e})_{1:i}))$ (9)
5.2 Experiments
1. Attention self-repair in large language models: We investigate whether the attention self-repair problem occurs in large language models and evaluate the effectiveness of TSG in amplifying the signal when gradients would otherwise have underestimated the importance of attention scores.
We first identify attention score vectors where multiple positions contribute significantly to the model output. We define position $j$ as having a significant contribution when:
$\frac{\partial z}{\partial s_{j}}s_{j}>\beta$
where $\beta$ is a threshold parameter.
For each identified attention vector, we measure the joint effect by ablating the two largest attention scores simultaneously versus the sum of their individual ablation effects. If the joint effect is larger than the sum of the individual effect, the attention scores exhibit OR-gate-like behaviour, which is equivalent to self-repair [26]. We quantify these effects by the change in weighted gradient sum $\sum_{k}\frac{\partial z}{\partial s_{k}}s_{k}$. We do this instead of computing the full forward pass to avoid other components compensating for the ablations.
Next, we compare TSG and standard gradient attribution scores for attention scores, focusing on examples with large self-repair effects. This comparison evaluates whether TSG correctly amplifies attribution for attention scores that standard gradients would underestimate due to self-repair.
2. Faithfulness of GIM: For feature attribution evaluation, we compare GIM with five gradient-based methods: GradientXInput (GxI) [5], Integrated gradients (IG) [5], DeepLIFT [27], TransformerLRP [9], and AttnLRP [28]. We did not evaluate perturbation-based methods because they are too computationally expensive for large language models on long inputs [29]. We computed comprehensiveness and sufficiency by replacing tokens with the token representing a space.
For circuit identification evaluation, we compare GIM with attribution patching (ATP) [29], ATP* [^2] We only implemented ATP*’s softmax gradient fix component. [15], and Integrated Gradients [28] by computing comprehensiveness and sufficiency at each layer separately, by replacing the representations by the layer-average representations.
All attribution scores follow the formula $(\bm{x}-\bm{\tilde{x}})\cdot\bm{\alpha}$, where $\bm{x}$ and $\bm{\tilde{x}}$ are original and counterfactual representations, and $\bm{\alpha}$ represents method-specific attributions. For the feature attribution evaluation, we used the token representing a space as the counterfactuals, while for the circuit identification evaluation, we used the layer-average representations as counterfactuals. We used a temperature of 2 for TSG, which we chose based on the results on Gemma-2 2B on the FEVER and HateXplain datasets.
3. Identifying the importance of each GIM modification: Our final experiment is an ablation study, where we estimate the importance of the three modifications in GIM. We estimate their importance by cumulatively adding modifications to standard GradientXInput and measuring the change in feature attribution faithfulness.
(a) BoolQ
(b) Fever
(c) Movie review
Figure 3: Comparison of circuit identification methods for LLAMA-3.2 1B (95% CI). The top row depicts comprehensiveness per layer ($\uparrow$). The bottom row depicts sufficiency ($\downarrow$).
6 Results
Self-repair is prevalent in attention mechanisms: Our analysis reveals that attention self-repair is common across large language models. Figure 2(a) demonstrates this phenomenon by comparing the joint effect of ablating the two largest attention scores simultaneously (y-axis) versus the sum of their individual ablation effects (x-axis). The consistent gap above the diagonal, especially prominent for orange-highlighted instances (top 5% strongest interactions), provides clear empirical evidence of self-repair: individual attention scores show disproportionately small effects when ablated separately. These results validate our theoretical analysis from Section 3.1 in practical model settings. We provide results for other models and datasets in Section 9.9.
TSG attributes more importance to attention scores during attention self-repair: Figure 2(b) compares standard gradient attribution (x-axis) with our Temperature-adjusted Softmax Gradients (y-axis) for the same attention scores. The orange dots represent attention vectors identified as having strong self-repair interactions in the left figure. TSG consistently attributes higher importance to these scores. This shows that TSG successfully addresses the gradient cancellation issue identified in our theoretical analysis by increasing attribution magnitude where standard gradients would otherwise assign low importance due to self-repair. We provide similar results for other models and datasets in Section 9.9.
GIM is more faithful than other gradient-based circuit identification methods: Figure 3 shows comprehensiveness and sufficiency metrics across layers for LLAMA-3.2 1B. GIM significantly outperforms other circuit identification methods in early layers, while Integrated Gradients demonstrates slight advantages in middle and late layers. Similar patterns appear across other models and datasets (Section 9.8).
For all the methods, the scores worsen in the top layers of the LLM. This could either be because they become less faithful or because the upper and lower limits of comprehensiveness and sufficiency change across layers [18]. Since we are interested in the relative differences between the methods, we leave it to future work to understand what causes the drop in the scores in top layers.
GIM is often more faithful than other gradient-based feature attribution methods: Table 1 compares GIM against five other gradient-based attribution methods across multiple datasets and model architectures. GIM drastically outperforms traditional methods like GradientXInput, Integrated Gradients, and DeepLIFT on nearly all evaluations. While AttnLRP [10] outperforms GIM on several dataset-model pairs, GIM achieves the highest comprehensiveness and sufficiency scores for the majority of combinations, establishing it as the most consistently faithful method overall.
Table 1: Comparison of gradient-based feature attribution methods.
Comprehensiveness $\uparrow$ Sufficiency $\downarrow$
Gemma LLAMA Qwen Gemma LLAMA Qwen
2B 9B 1B 3B 8B 1.5B 3B 2B 9B 1B 3B 8B 1.5B 3B
BoolQ GxI 0.09 0.00 0.18 0.45 0.39 0.27 0.57 0.60 0.43 0.71 0.39 0.67 0.54 0.43
IG 0.51 0.20 0.20 0.52 0.40 0.00 0.48 0.11 0.12 0.43 0.34 0.58 0.63 0.54
DeepLIFT 0.41 0.15 0.26 0.36 0.31 0.20 0.34 0.37 0.37 0.58 0.55 0.74 0.67 0.77
TransformerLRP 0.32 0.09 0.57 0.55 0.17 0.26 0.27 0.37 0.07 0.30 0.30 0.68 0.49 0.64
AttnLRP - - 0.66 0.70 0.61 0.67 0.65 - - 0.25 0.17 0.50 0.15 0.34
GIM 0.59 0.43 0.69 0.72 0.60 0.68 0.61 0.03 0.04 0.22 0.10 0.49 0.09 0.23
FEVER GxI 0.03 0.06 0.39 0.53 0.46 0.43 0.47 0.47 0.40 0.56 0.63 0.71 0.59 0.73
IG 0.47 0.32 0.28 0.51 0.50 0.03 0.54 0.02 0.20 0.51 0.59 0.62 0.67 0.61
DeepLIFT 0.12 0.20 0.33 0.47 0.53 0.31 0.43 0.43 0.36 0.58 0.65 0.76 0.64 0.76
TransformerLRP 0.17 0.22 0.54 0.68 0.21 0.30 0.49 0.44 0.21 0.36 0.46 0.75 0.63 0.69
AttnLRP - - 0.60 0.75 0.65 0.61 0.68 - - 0.32 0.44 0.52 0.23 0.49
GIM 0.42 0.39 0.62 0.75 0.69 0.51 0.68 -0.01 0.09 0.26 0.39 0.51 0.31 0.41
HateXplain GxI 0.07 0.08 0.53 0.38 0.44 0.49 0.53 0.59 0.42 0.60 0.67 0.73 0.63 0.63
IG 0.62 0.37 0.49 0.64 0.44 0.00 0.46 0.35 0.32 0.58 0.56 0.61 0.64 0.72
DeepLIFT 0.36 0.29 0.55 0.36 0.41 0.29 0.42 0.52 0.45 0.59 0.70 0.73 0.74 0.73
TransformerLRP 0.54 0.14 0.51 0.58 0.33 0.20 0.30 0.38 0.24 0.59 0.57 0.67 0.66 0.71
AttnLRP - - 0.68 0.70 0.60 0.85 0.59 - - 0.30 0.47 0.42 0.33 0.71
GIM 0.64 0.48 0.68 0.64 0.65 0.80 0.63 0.26 0.14 0.26 0.46 0.47 0.35 0.51
Movie GxI 0.15 0.04 0.34 0.51 0.34 0.42 0.52 0.61 0.52 0.61 0.51 0.71 0.68 0.76
IG 0.60 0.38 0.28 0.69 0.41 0.00 0.58 0.17 0.13 0.49 0.36 0.52 0.71 0.56
DeepLIFT 0.44 0.22 0.31 0.62 0.48 0.49 0.35 0.35 0.48 0.60 0.49 0.71 0.56 0.77
TransformerLRP 0.48 0.26 0.61 0.70 0.22 0.23 0.49 0.27 0.13 0.35 0.40 0.83 0.74 0.67
AttnLRP - - 0.69 0.76 0.72 0.78 0.77 - - 0.25 0.30 0.48 0.19 0.46
GIM 0.60 0.44 0.71 0.78 0.71 0.83 0.76 0.11 0.07 0.19 0.30 0.48 0.17 0.31
SciFact GxI 0.09 0.05 0.27 0.42 0.40 0.34 0.57 0.55 0.39 0.57 0.59 0.72 0.56 0.63
IG 0.58 0.32 0.27 0.46 0.50 0.01 0.65 0.18 0.20 0.46 0.56 0.57 0.69 0.54
DeepLIFT 0.28 0.27 0.28 0.43 0.50 0.25 0.62 0.47 0.37 0.58 0.61 0.75 0.69 0.65
TransformerLRP 0.30 0.22 0.56 0.56 0.15 0.39 0.51 0.41 0.22 0.33 0.43 0.76 0.50 0.67
AttnLRP - - 0.64 0.71 0.65 0.67 0.74 - - 0.30 0.35 0.49 0.22 0.40
GIM 0.57 0.40 0.67 0.74 0.67 0.62 0.74 0.04 0.08 0.22 0.31 0.49 0.26 0.38
Twitter GxI 0.40 0.27 0.41 0.48 0.45 0.48 0.73 0.51 0.37 0.64 0.75 0.77 0.73 0.64
IG 0.53 0.34 0.41 0.70 0.55 0.40 0.67 0.36 0.33 0.62 0.59 0.66 0.71 0.66
DeepLIFT 0.44 0.23 0.42 0.65 0.54 0.51 0.75 0.54 0.44 0.62 0.67 0.75 0.75 0.67
TransformerLRP 0.43 0.26 0.58 0.62 0.34 0.41 0.50 0.43 0.38 0.49 0.66 0.76 0.73 0.72
AttnLRP - - 0.64 0.73 0.68 0.75 0.66 - - 0.42 0.61 0.57 0.48 0.66
GIM 0.52 0.38 0.68 0.72 0.69 0.76 0.72 0.34 0.22 0.41 0.61 0.56 0.55 0.58
All modifications contribute to GIM’s faithfulness improvements: Figure 4 compares how adding different GIM modifications to GradientXInput improves feature attribution faithfulness. Each point represents the average improvement for a dataset compared to the GradientXInput baseline. Blue boxplots show feature attribution performance with LayerNorm freeze + Gradient Normalization, while orange boxplots include TSG as well (full GIM). The results demonstrate that LayerNorm freeze + Gradient Normalization substantially improve both attribution metrics across all models. Adding TSG further enhances attribution quality for most models, with particularly strong improvements for Gemma-2 2B and LLAMA models. However, TSG appears less beneficial for Qwen-2.5 1.5B, where it occasionally reduces attribution faithfulness. Qwen-2.5 1.5B had a hundred-fold larger gradients than the other models, which perhaps is problematic for TSG (see Figure 11). Comprehensive feature attribution results for all modification combinations are available in Table 2.
(a) Comprehensiveness $\uparrow$
(b) Sufficiency $\downarrow$
Figure 4: Relative improvement of adding modifications to GradientXInput. We compare adding layernorm freeze and grad norm with adding also TSG. Each point represents the average statistics of a dataset. The y-axis shows the improvements over GradientXInput.
7 Discussion
When to amplify signals during self-repair: Our temperature-adjusted softmax gradients (TSG) amplify attribution signals when attention self-repair occurs, but is this amplification justified? We argue that amplification is appropriate for components that were actively contributing to the output before any perturbation, even when self-repair dampens the observed output change. The “firing squad” analogy from Pearl [30] illustrates this well: when two soldiers simultaneously fire at a prisoner (where either shot alone would be fatal), both bear full responsibility despite the redundancy. Even though removing one shooter’s bullet would not change the outcome (a form of self-repair), we should still attribute causal responsibility to both shooters.
Crucially, we should not attribute importance to components that only begin to contribute after perturbation occurs, regardless of whether this redistribution happens immediately (as in attention self-repair) or through backup mechanisms (as in the hydra effect [8]). Consider a historical counterfactual: “What killed Archduke Franz Ferdinand?” If we ablate the assassination in Sarajevo, and in this counterfactual he later dies of cancer, we would not say he died from both being shot and cancer. When a component only contributes after ablation of another component, that compensating component should not receive attribution in the unperturbed state.
Could GIM improve model training? Since GIM’s gradient modifications more accurately attribute importance to model components during interpretation, we speculate that they might also improve model training itself. By implementing these modifications during backpropagation, we might achieve more precise parameter updates that account for feature interactions and self-repair mechanisms. We suggest that future work empirically evaluate whether models trained with GIM-modified gradients demonstrate faster convergence or better performance compared to standard training approaches.
Limitations: While our experiments demonstrate that TSG increases both backpropagation signals during self-repair and overall attribution faithfulness, we cannot conclusively establish a causal link between these observations. Temperature adjustment affects backpropagation in multiple ways beyond mitigating self-repair, and these additional effects may contribute to the improved faithfulness we observe.
Furthermore, we did not find an explanation for why TSG improved faithfulness across all models except for Qwen-2.5 1.5B. We observed that its gradient was a hundred-fold larger than the other models’, but could not provide any definite evidence that this caused the decreased faithfulness.
8 Conclusion
We identified attention self-repair, a phenomenon where softmax normalization redistributes weights during perturbation, causing interpretability methods to underestimate component importance. Our proposed method, Gradient Interaction Modifications (GIM), addresses this through temperature-adjusted softmax gradients (TSG), while also incorporating layer normalization freeze and gradient normalization to handle other issues in transformer interpretability. Our evaluations across multiple models and datasets show that GIM consistently outperforms existing circuit identification and feature attribution methods in faithfulness. By accounting for self-repair during backpropagation, GIM better quantifies the importance of model components, advancing our understanding of neural networks by improving our ability to interpret them.
Acknowledgments and Disclosure of Funding
Thank you, Casper L. Christensen, for the enthusiastic discussions and for proposing and testing ideas for how to improve GIM. Thank you, Jakob Havtorn, Lasse Borgholt, and Søren Hauberg, for discussions and advice.
This research was partially funded by the Innovation Fund Denmark via the Industrial Ph.D. Program (grant no. 2050-00040B) and Academy of Finland (grant no. 322653).